# Perturbative soft graviton theorems in de Sitter spacetime

### v2

@@ 4.2 Mode Solutions
where $\Box=\partial^{\gamma}\partial_{\gamma}$ is the flat space de-Alembertian operator. The above expression is the flat space linearized equation with perturbative corrections up to $\mathcal{O}(l^{-2})$. The de Donder like gauge condition takes the following form in the stereographic coordinates:
$\left(1+\dfrac{x^{2}}{2l^{2}}\right)\left(\partial^{\alpha}h_{\nu\alpha}-\dfrac{1}{2}\eta^{\alpha\beta}\partial_{\nu}h_{\alpha\beta}\right)-\dfrac{x^{\alpha}h_{\nu\alpha}}{l^{2}}=0,$ (34)
which for $l\to\infty$ reduces to the flat space de Donder gauge. To determine mode solutions of the linearized field equation, we use the following ansatz:
$f^{h}_{\mu\nu}(x,p)=\dfrac{1}{\sqrt{2E_{p}}}\varepsilon^{h}_{\mu\nu}(p)\left(1+\dfrac{\mathcal{F}(x)}{l^{2}}\right)e^{ip\cdot x},$ (35)
where $\varepsilon_{\mu\nu}$ is the polarization tensor, $E_{p}$ is the zeroth component of $p^{\mu}$, “$h$” indicates the helicity of the graviton [^2] The helicity index “$h$” should not be confused with the field $h_{\mu\nu}$. , and $\mathcal{F}(x)$ is an arbitrary function that can be obtained by substituting the ansatz in (33) and solving it for $\mathcal{F}(x)$. We obtained:
$\mathcal{F}(x)=\dfrac{x^{2}}{2},\ p^{2}=\frac{6}{l^{2}}.$ (36)
For $l\to\infty$, eq. (35) reduces to the plane wave solution for the flat space linearized field equation. Given the mode solutions, we note that the gauge condition (34) implies $p^{\alpha}\varepsilon^{h}_{\alpha\nu}=0$, and hence $\varepsilon^{\mu}_{\mu}=0$.
@@
where $a^{h}_{p}$ ($a^{h\dagger}_{p}$) is the annihilation (creation) operator which acts on the asymptotic states to annihilate (create) a graviton of momentum $p$ and helicity “$h$”.
@@ 4.3 LSZ Formula for Gravitons
Since the scattering processes are confined to a compact region $R$ within the static patch, the boundary of $R$ can be achieved in a finite time $T$ which we assume to be larger than the interaction time scale. Therefore,
[…]
@@
where the operator $\left(\nabla_{x}^{2}-2/l^{2}\right)$ acts on the graviton propagator (or Green’s function) $G_{\mu\nu\alpha\beta}(x-x^{\prime})$ as follows:
@@ 5 Perturbative $\mathcal{S}$-matrix and LSZ Formula
In this section, we define a perturbative $\mathcal{S}$-matrix in the de Sitter space and deduce curved spacetime generalization of the LSZ formula. The proper description of the $\mathcal{S}$-matrix requires well-defined asymptotic states at the boundaries of the spacetime. Since the largest observable region in de Sitter space is the static patch, which is bounded by cosmological horizons, we can not define asymptotic states at the boundaries of the de Sitter space. In such situations, one can treat cosmological horizons as physical boundaries [^3] In Teitelboim_2001 ; Teitelboim_2003 , the cosmological horizons are considered as physical boundaries to obtain the thermodynamics of a rotating black hole. of the spacetime to define the asymptotic states. In our analysis, the scattering processes are confined to the compact region $R$ inside the static patch. Thus, the asymptotic states can be defined at the boundaries of $R$. All lengths in our calculations are very small compared to the de Sitter radius $l$, i.e. $|x|<<l$. Also, the time $T$ at the boundaries of $R$ is considered to be larger than the interaction time scale.
@@
6 Soft Graviton Theorems
@@
In this section, we compute perturbative corrections ($\mathcal{O}(l^{-1})$ and $\mathcal{O}(l^{-2})$ corrections) to the flat space soft graviton theorem using the LSZ formula discussed in the previous section, and deduce leading, subleading and sub-subleading corrections. We consider a scattering of $n$ scalars with momenta $p_{i}$ followed by an emission of a soft graviton with momentum $k$, as shown in fig. (2). The graviton is attached to an external line by three-point minimal coupling. The graviton attached to an internal line or to higher-point couplings are not relevant to our computation. We consider all external lines to be outgoing for the ease of the calculations. The $\mathcal{S}$-matrix corresponds to the scattering process depicted in fig. (2) is described as
@@
Figure 2: Scattering of scalars followed by an emission of a soft graviton.
@@
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}\kappa_{i}\int d^{4}z_{i}\sqrt{-g(z_{i})}g^{*}_{p_{i}}(z_{i})(-i)(\nabla^{2}_{z_{i}}-m^{2})\\ \cdot\int d^{4}y\sqrt{-g(y)}\mathcal{P}^{\mu\nu\alpha\beta}f^{*}_{\mu\nu}(y,k)(-i)(\nabla_{y}^{2}-2/l^{2})\\ \cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}\sqrt{-g(z_{j})}g^{*}_{p_{j}}(z_{j})(-i)(\nabla^{2}_{z_{j}}-m^{2})\int d^{4}z\sqrt{-g(z)}\int d^{4}z^{\prime}\sqrt{-g(z^{\prime})}\\ \cdot\bigg{\langle}\mathcal{T}\phi(z_{i})h_{\alpha\beta}(y)h^{\rho\sigma}(z^{\prime})\bigg{[}\partial^{\prime}_{\rho}\phi(z^{\prime})\partial^{\prime}_{\sigma}\phi(z^{\prime})-\frac{1}{2}g_{\rho\sigma}\partial^{\prime}_{\tau}\phi(z^{\prime})\partial^{\prime\tau}\phi(z^{\prime})+\frac{1}{2}g_{\rho\sigma}m^{2}\phi^{2}(z^{\prime})\bigg{]}G[\phi(z)...\phi(z_{j})]\bigg{\rangle},$ (49)
@@
where $G[\phi(z)...\phi(z_{j})]$ is the $n$-point correlation function for scalars, and $\kappa_{i}$ is the scalar-graviton coupling constant. The sum is taken over all the scalars which can emit a soft graviton to take into account the contributions from all the external legs. Now, we perform the Wick contraction among the fields and generate the respective propagators.
@@
$\bigg{\langle}\mathcal{T}\phi(z_{i})h_{\alpha\beta}(y)h^{\rho\sigma}(z^{\prime})\bigg{[}\partial^{\prime}_{\rho}\phi(z^{\prime})\partial^{\prime}_{\sigma}\phi(z^{\prime})-\frac{1}{2}g_{\rho\sigma}\partial^{\prime}_{\tau}\phi(z^{\prime})\partial^{\prime\tau}\phi(z^{\prime})+\frac{1}{2}g_{\rho\sigma}m^{2}\phi^{2}(z^{\prime})\bigg{]}G[\phi(z)...\phi(z_{j})]\bigg{\rangle}\\ =\bigg{\langle}G_{\alpha\beta}^{\rho\sigma}(y-z^{\prime})\big{\{}2\partial^{\prime}_{\rho}D(z_{i},z^{\prime})\partial^{\prime}_{\sigma}D(z^{\prime},z)-g_{\rho\sigma}\partial^{\prime}_{\tau}D(z_{i},z^{\prime})\partial^{\prime\tau}D(z^{\prime},z)\\ +g_{\rho\sigma}m^{2}D(z_{i},z^{\prime})D(z^{\prime},z)\big{\}}G[\{\phi(z_{j})\}]\bigg{\rangle},$ (50)
@@
where $G^{\rho\sigma}_{\alpha\beta}(x-y)$ and $D(x,y)$ are the graviton and scalar propagators, respectively. Since the propagators are the Green’s functions, the operators in the above expression, $(\nabla_{z_{i}}^{2}-m^{2})$ and $(\nabla_{y}^{2}-2/l^{2})$, will act on the respective propagator to yield the Dirac delta functions as follows:
@@
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}\kappa_{i}\int d^{4}z_{i}\sqrt{-g(z_{i})}g^{*}_{p_{i}}(z_{i})(-i)\int d^{4}y\sqrt{-g(y)}\mathcal{P}^{\mu\nu\alpha\beta}f^{*}_{\mu\nu}(y,k)(-i)\\ \cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}\sqrt{-g(z_{j})}g^{*}_{p_{j}}(z_{j})(-i)(\nabla^{2}_{z_{j}}-m^{2})\int d^{4}z\sqrt{-g(z)}\int d^{4}z^{\prime}\sqrt{-g(z^{\prime})}\\ \cdot\bigg{\langle}\mathcal{P}_{\alpha\beta}^{\rho\sigma}\dfrac{i\delta^{(4)}(y-z^{\prime})}{\sqrt{-g(y)}}\bigg{\{}2i\partial^{\prime}_{\rho}\dfrac{\delta^{(4)}(z_{i}-z^{\prime})}{\sqrt{-g(z_{i})}}\partial^{\prime}_{\sigma}D(z^{\prime},z)-ig_{\rho\sigma}\partial^{\prime}_{\tau}\dfrac{\delta^{(4)}(z_{i}-z^{\prime})}{\sqrt{-g(z_{i})}}\partial^{\prime\tau}D(z^{\prime},z)\\ +ig_{\rho\sigma}m^{2}\dfrac{\delta^{(4)}(z_{i}-z^{\prime})}{\sqrt{-g(z_{i})}}D(z^{\prime},z)\bigg{\}}G[\{\phi(z_{j})\}]\bigg{\rangle}.$ (51)
@@
Let us now evaluate integrations in $y$ and $z_{i}$. First, we perform the $y$-integral which is very trivial due to the presence of the delta function $\delta^{(4)}(y-z^{\prime})$. After performing it, the coordinate $y$ in the above expression will simply be replaced by $z^{\prime}$. Similarly, we can perform the $z_{i}$-integral on another delta function $\delta^{(4)}(z_{i}-z^{\prime})$. Since the coordinates $z_{i}$ and $z^{\prime}$ are independent, we move the $z_{i}$-integral inside the derivative operator $\partial^{\prime}_{\alpha}$, which acts on the delta function in the above expression, and then perform the integration. We obtain
@@
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},k)=\int d^{4}z\sqrt{-g(z)}\int d^{4}z^{\prime}\sqrt{-g(z^{\prime})}\mathcal{P}^{\mu\nu\alpha\beta}\mathcal{P}_{\alpha\beta}^{\rho\sigma}f^{*}_{\mu\nu}(z^{\prime},k)\\ \cdot\sum_{i=1}^{n}\kappa_{i}\bigg{\{}2\partial^{\prime}_{\rho}g^{*}_{p_{i}}(z^{\prime})\partial^{\prime}_{\sigma}D(z^{\prime},z)-g_{\rho\sigma}\partial^{\prime}_{\tau}g^{*}_{p_{i}}(z^{\prime})\partial^{\prime\tau}D(z^{\prime},z)+g_{\rho\sigma}m^{2}g^{*}_{p_{i}}(z^{\prime})D(z^{\prime},z)\bigg{\}}\\ \cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}\sqrt{-g(z_{j})}g^{*}_{p_{j}}(z_{j})(-i)(\nabla^{2}_{z_{j}}-m^{2})\big{\langle}G[\{\phi(z_{j})\}]\big{\rangle},$ (52)
@@
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}\kappa_{i}\int d^{4}zd^{4}z^{\prime}\left(1-\dfrac{z^{2}}{l^{2}}-\dfrac{z^{\prime 2}}{l^{2}}\right)f^{*\rho\sigma}(z^{\prime},k)\\ \cdot\bigg{\{}2\partial^{\prime}_{\rho}g^{*}_{p_{i}}(z^{\prime})\partial^{\prime}_{\sigma}D(z^{\prime},z)-g_{\rho\sigma}\partial^{\prime}_{\tau}g^{*}_{p_{i}}(z^{\prime})\partial^{\prime\tau}D(z^{\prime},z)+g_{\rho\sigma}m^{2}g^{*}_{p_{i}}(z^{\prime})D(z^{\prime},z)\bigg{\}}\\ \int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}\sqrt{-g(z_{j})}g^{*}_{p_{j}}(z_{j})(-i)(\nabla^{2}_{z_{j}}-m^{2})\big{\langle}G[\{\phi(z_{j})\}]\big{\rangle}.$ (53)
Let us now do the integration by parts in $z^{\prime}$ in the above expression, and substitute the following identities: $\partial^{\prime}_{\rho}f^{*\rho\sigma}=\dfrac{z^{\prime}_{\rho}}{l^{2}}f^{*\rho\sigma}$, $\partial^{\prime}_{\rho}\sqrt{-g(z^{\prime})}=-2\dfrac{z^{\prime}_{\rho}}{l^{2}}$, $g_{\rho\sigma}\varepsilon^{*\rho\sigma}=0$. We obtain
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},k)=-\sum_{i=1}^{n}2\kappa_{i}\int d^{4}zd^{4}z^{\prime}\left(1-\dfrac{z^{2}}{l^{2}}\right)f^{*\rho\sigma}(z^{\prime},k)g^{*}_{p_{i}}(z^{\prime})\\ \cdot\bigg{\{}\left(1-\dfrac{z^{\prime 2}}{l^{2}}\right)\partial^{\prime}_{\rho}\partial^{\prime}_{\sigma}D(z^{\prime},z)-\dfrac{z^{\prime}_{\rho}}{l^{2}}\partial^{\prime}_{\sigma}D(z^{\prime},z)\bigg{\}}\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}\sqrt{-g(z_{j})}g^{*}_{p_{j}}(z_{j})(-i)(\nabla^{2}_{z_{j}}-m^{2})\big{\langle}G[\{\phi(z_{j})\}]\big{\rangle}.$ (54)
The above expression can be put in the following form after substituting all essential elements and taking the soft limit, i.e., $\omega\to 0$ (or $k\to 0$).
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},\omega\hat{k})=\bigg{(}A^{(0)}(\{p_{i}\},\omega\hat{k})+\dfrac{1}{l}A^{(1)}(\{p_{i}\},\omega\hat{k})+\dfrac{1}{l^{2}}A^{(2)}(\{p_{i}\},\omega\hat{k})\bigg{)}\\ \cdot\bigg{(}\Gamma_{n}^{(0)}(\{p_{i}\})+\dfrac{1}{l}\Gamma_{n}^{(1)}(\{p_{i}\})+\dfrac{1}{l^{2}}\Gamma_{n}^{(2)}(\{p_{i}\})\bigg{)},$ (55)

### v3

@@ 4.2 Mode solutions
where $\Box=\partial^{\gamma}\partial_{\gamma}$ is the flat space d’Alembertian operator. The above expression is the flat space linearized equation with perturbative corrections up to $\mathcal{O}(l^{-2})$.
Now, we determine the mode solutions for the linearized field equation by considering the ansatz as follows:
$f^{h}_{\mu\nu}(x,p)=\dfrac{1}{\sqrt{2E_{p}}}\left(\varepsilon^{h}_{\mu\nu}(p)+\varepsilon^{h}_{\mu\alpha}\dfrac{\mathcal{F}^{\alpha}_{\nu}(x)}{l^{2}}+\varepsilon^{h}_{\alpha\nu}\dfrac{\mathcal{F}_{\mu}^{\alpha}(x)}{l^{2}}+\varepsilon^{h}_{\alpha\beta}\dfrac{\mathcal{G}^{\alpha\beta}_{\ \ \ \mu\nu}(x)}{l^{2}}\right)e^{ip\cdot x},$ (39)
where $\varepsilon_{\mu\nu}$ is the polarization tensor, $E_{p}$ is the zeroth component of $p^{\mu}$, “hh” indicates the helicity of the graviton [^3] The helicity index “hh” should not be confused with the trace of $h_{\mu\nu}$. , and $\mathcal{F}_{\alpha\beta}$ and $\mathcal{G}_{\alpha\beta\mu\nu}$ are arbitrary tensors on the flat space that can be obtained by substituting the ansatz (39) into (38) and solving it for $\mathcal{F}_{\alpha\beta}$ and $\mathcal{G}_{\alpha\beta\mu\nu}$. We obtained:
$\mathcal{F}_{\alpha\beta}(x)=\frac{1}{2}x_{\alpha}x_{\beta}-\frac{1}{8}\eta_{\alpha\beta}x^{2},$ (40)
$\mathcal{G}_{\alpha\beta\mu\nu}=-\frac{1}{4}x_{\alpha}x_{\beta}\eta_{\mu\nu},\ \ p^{2}=\frac{2}{l^{2}};$ (41)
where $\varepsilon^{\alpha}_{\alpha}=0$ and $p^{\alpha}\varepsilon_{\alpha\beta}=\mathcal{O}(l^{-2})$ have been used to obtain the solution, which are consistent with the transverse-traceless gauge (37). For $l\to\infty$, the mode solution reduces to the plane wave solution for the flat space linearized field equation.
@@
where $a^{h}_{p}$ ($a^{h\dagger}_{p}$) is the annihilation (creation) operator which acts on the asymptotic states to annihilate (create) a graviton of momentum $p$ and helicity “hh”.
@@ 4.3 LSZ formula for gravitons
Since the scattering processes are confined to a compact region $R$ inside the static patch, the particles can reach the boundaries of $R$ in a finite time $T$ which is much larger than the interaction time scale. Therefore,
$a_{p}^{h}(T)-a_{p}^{h}(-T)=\int_{-T}^{T}dt\ \partial_{t}a_{p}^{h}=\int_{-T}^{T}dt\ \partial_{t}\left(\mathcal{P}^{\mu\nu\alpha\beta}f^{h}_{\alpha\beta}(x,p),h_{\mu\nu}(x,p^{\prime})\right)\\ =-i\int dtd^{3}x\ e^{-\epsilon\frac{|\overrightarrow{x}|}{R}}\partial_{t}\left(\sqrt{-g}\mathcal{P}^{\mu\nu\alpha\beta}\ \big(f^{*h}_{\alpha\beta}(x,p)\nabla^{t}h_{\mu\nu}(x,p^{\prime})-h_{\alpha\beta}(x,p^{\prime})\nabla^{t}f^{*h}_{\mu\nu}(x,p)\big)\right)\\ =-i\int d^{4}x\ e^{-\epsilon\frac{|\overrightarrow{x}|}{R}}\partial_{\rho}\left(\sqrt{-g}\mathcal{P}^{\mu\nu\alpha\beta}\ \big(f^{*h}_{\alpha\beta}(x,p)\nabla^{\rho}h_{\mu\nu}(x,p^{\prime})-h_{\alpha\beta}(x,p^{\prime})\nabla^{\rho}f^{*h}_{\mu\nu}(x,p)\big)\right)\\ =-i\int d^{4}xe^{-\epsilon\frac{|\overrightarrow{x}|}{R}}\sqrt{-g}\ \nabla_{\rho}\left(\mathcal{P}^{\mu\nu\alpha\beta}\ \big(f^{*h}_{\alpha\beta}(x,p)\nabla^{\rho}h_{\mu\nu}(x,p^{\prime})-h_{\alpha\beta}(x,p^{\prime})\nabla^{\rho}f^{*h}_{\mu\nu}(x,p)\big)\right),$
@@
where the operator $\left(\nabla^{2}-2/l^{2}\right)$ acts on the graviton propagator (or Green’s function) $G_{\mu\nu\alpha\beta}(x-x^{\prime})$ as follows:
@@ 5 Perturbative $\mathcal{S}$-matrix and LSZ formula
In this section, we define a perturbative $\mathcal{S}$-matrix in the de Sitter space and deduce curved spacetime generalization of the LSZ formula. The proper description of the $\mathcal{S}$-matrix requires well-defined asymptotic states at the boundaries of the spacetime. Since the largest observable region in de Sitter space is the static patch, which is bounded by cosmological horizons, we can not define asymptotic states at the boundaries of the de Sitter space. In such situations, one can treat cosmological horizons as physical boundaries [^4] In Teitelboim_2001 ; Teitelboim_2003 , the cosmological horizons are considered as physical boundaries to study the thermodynamics of a rotating black hole. of the spacetime to define the asymptotic states. In our analysis, since the scattering processes are confined to the compact region inside the static patch, the asymptotic states are defined at the early- and late-time Cauchy surfaces which act as Hilbert spaces of the incoming and outgoing particles. The particles can reach the Cauchy surface in a finite time $T$, which we assume to be much larger than the interaction time scale. All lengths in our calculations are very small compared to the de Sitter radius $l$, i.e., $|x|<<l$.
@@
6 Perturbative corrections to flat space soft graviton theorem
In this section, we compute perturbative corrections ($\mathcal{O}(l^{-1})$ and $\mathcal{O}(l^{-2})$ corrections) to the flat space soft graviton theorem using the LSZ formula discussed in the previous section, and deduce leading, subleading, and sub-subleading corrections. The action for a massive scalar field minimally coupled to the Einstein gravity is written as
$S=-\int d^{4}x\sqrt{-g}\bigg[\frac{2}{\kappa^{2}}(\mathcal{R}-2\Lambda)+\frac{1}{2}g^{\mu\nu}\nabla_{\mu}\phi\nabla_{\nu}\phi-\frac{1}{2}m^{2}\phi^{2}+V(\phi)\bigg],$ (54)
where $V(\phi)$ is a scalar potential, $g_{\mu\nu}$ is a general spacetime metric, and $\nabla_{\mu}$ is the covariant derivative compatible with $g_{\mu\nu}$. In the weak field expansion, $g_{\mu\nu}=\bar{g}_{\mu\nu}+\kappa h_{\mu\nu}$, [^5] Bar notations are used for the de Sitter space to distinguish it from the general spacetime, which will be dropped later. the leading order terms are
$\mathcal{L}_{gravity}$ $=-\frac{2}{\kappa^{2}}\sqrt{-g}(\mathcal{R}-2\Lambda)$
$=\frac{1}{2}\bar{\nabla}_{\sigma}h_{\mu\nu}\bar{\nabla}^{\sigma}h^{\mu\nu}-\frac{1}{2}\bar{\nabla}_{\sigma}h\bar{\nabla}^{\sigma}h-\bar{\nabla}_{\mu}h_{\nu\sigma}\bar{\nabla}^{\sigma}h^{\mu\nu}+\bar{\nabla}_{\mu}h^{\mu\nu}\bar{\nabla}_{\nu}h-\frac{3}{l^{2}}\bigg(h_{\mu\nu}h^{\mu\nu}-\frac{1}{2}h^{2}\bigg)+\cdots,$ (55)
$\mathcal{L}_{scalar}$ $=-\sqrt{-g}\bigg[\frac{1}{2}g^{\mu\nu}\nabla_{\mu}\phi\nabla_{\nu}\phi-\frac{1}{2}m^{2}\phi^{2}+V(\phi)\bigg]$
$=-\frac{1}{2}\bar{\nabla}_{\sigma}\phi\bar{\nabla}^{\sigma}\phi+\frac{1}{2}m^{2}\phi^{2}-V(\phi)+\frac{\kappa}{2}h^{\mu\nu}\bigg[\bar{\nabla}_{\mu}\phi\bar{\nabla}_{\nu}\phi-\frac{1}{2}\bar{g}_{\mu\nu}\big(\bar{\nabla}_{\sigma}\phi\bar{\nabla}^{\sigma}\phi-m^{2}\phi^{2}\big)\bigg]+\cdots,$ (56)
where the total derivative terms and the constant terms have been dropped in the Lagrangian.
@@ 6 Perturbative corrections to flat space soft graviton theorem
Let us consider a scattering of $n$ scalars with momenta $p_{i}$ followed by an emission of a soft graviton with momentum $k$, as shown in Fig. (2). The graviton is attached to an external line by three-point minimal coupling. The graviton attached to an internal line, i.e., Fig. (3), will be analyzed in subsection (6.3). We consider all external lines to be outgoing for the ease of the calculations. The $\mathcal{S}$-matrix corresponds to the scattering process depicted in Fig. (2) is described as follows: [^6] Hereafter, the bar notations for the de Sitter space have been dropped.
@@
Figure 2: Scattering of scalars with momenta $p_{i}$ followed by an emission of a soft graviton with momentum $k$.
@@
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}\frac{\kappa_{i}}{2}\int d^{4}z_{i}\sqrt{-g(z_{i})}g^{*}_{p_{i}}(z_{i})(-i)(\nabla^{2}_{z_{i}}-m^{2})\\ \cdot\int d^{4}y\sqrt{-g(y)}\mathcal{P}^{\mu\nu\alpha\beta}f^{*}_{\mu\nu}(y,k)(-i)(\nabla_{y}^{2}-2/l^{2})\\ \cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}\sqrt{-g(z_{j})}g^{*}_{p_{j}}(z_{j})(-i)(\nabla^{2}_{z_{j}}-m^{2})\int d^{4}z\sqrt{-g(z)}\int d^{4}z^{\prime}\sqrt{-g(z^{\prime})}\\ \cdot\bigg\langle\mathcal{T}\phi(z_{i})h_{\alpha\beta}(y)h^{\rho\sigma}(z^{\prime})\bigg[\nabla^{\prime}_{\rho}\phi(z^{\prime})\nabla^{\prime}_{\sigma}\phi(z^{\prime})-\frac{1}{2}g_{\rho\sigma}\nabla^{\prime}_{\tau}\phi(z^{\prime})\nabla^{\prime\tau}\phi(z^{\prime})+\frac{1}{2}g_{\rho\sigma}m^{2}\phi^{2}(z^{\prime})\bigg]G[\phi(z)...\phi(z_{j})]\bigg\rangle,$ (57)
@@
where $G[\phi(z)...\phi(z_{j})]$ is the $n$-point correlation function for scalars which receives contributions from any arbitrary interaction vertices, and $\kappa_{i}$ is the scalar-graviton coupling constant. The sum is taken over all the scalars which can emit a soft graviton to take into account the contributions from all the external legs. Now, we perform the Wick contractions among the fields and generate the respective propagators as follows:
@@
$\bigg\langle\mathcal{T}\phi(z_{i})h_{\alpha\beta}(y)h^{\rho\sigma}(z^{\prime})\bigg[\nabla^{\prime}_{\rho}\phi(z^{\prime})\nabla^{\prime}_{\sigma}\phi(z^{\prime})-\frac{1}{2}g_{\rho\sigma}\nabla^{\prime}_{\tau}\phi(z^{\prime})\nabla^{\prime\tau}\phi(z^{\prime})+\frac{1}{2}g_{\rho\sigma}m^{2}\phi^{2}(z^{\prime})\bigg]G[\phi(z)...\phi(z_{j})]\bigg\rangle\\ =\bigg\langle G_{\alpha\beta}^{\rho\sigma}(y-z^{\prime})\big\{2\nabla^{\prime}_{\rho}D(z_{i},z^{\prime})\nabla^{\prime}_{\sigma}D(z^{\prime},z)-g_{\rho\sigma}\nabla^{\prime}_{\tau}D(z_{i},z^{\prime})\nabla^{\prime\tau}D(z^{\prime},z)\\ +g_{\rho\sigma}m^{2}D(z_{i},z^{\prime})D(z^{\prime},z)\big\}G[\{\phi(z_{j})\}]\bigg\rangle,$ (58)
@@
where $G^{\rho\sigma}_{\alpha\beta}(x-y)$ and $D(x,y)$ are graviton and scalar propagators, respectively. Since the propagators are the Green’s functions, the equation-of-motion operators in (57), i.e., $(\nabla_{z_{i}}^{2}-m^{2})$ and $(\nabla_{y}^{2}-2/l^{2})$, will act on the respective propagator and produce the delta functions.
@@
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}\frac{\kappa_{i}}{2}\int d^{4}z_{i}\sqrt{-g(z_{i})}g^{*}_{p_{i}}(z_{i})(-i)\int d^{4}y\sqrt{-g(y)}\mathcal{P}^{\mu\nu\alpha\beta}f^{*}_{\mu\nu}(y,k)(-i)\\ \cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}\sqrt{-g(z_{j})}g^{*}_{p_{j}}(z_{j})(-i)(\nabla^{2}_{z_{j}}-m^{2})\int d^{4}z\sqrt{-g(z)}\int d^{4}z^{\prime}\sqrt{-g(z^{\prime})}\\ \cdot\bigg\langle\mathcal{P}_{\alpha\beta}^{\rho\sigma}\dfrac{i\delta^{(4)}(y-z^{\prime})}{\sqrt{-g(y)}}\bigg\{2i\partial^{\prime}_{\rho}\dfrac{\delta^{(4)}(z_{i}-z^{\prime})}{\sqrt{-g(z_{i})}}\partial^{\prime}_{\sigma}D(z^{\prime},z)-ig_{\rho\sigma}g^{\tau\lambda}\partial^{\prime}_{\tau}\dfrac{\delta^{(4)}(z_{i}-z^{\prime})}{\sqrt{-g(z_{i})}}\partial^{\prime}_{\lambda}D(z^{\prime},z)\\ +ig_{\rho\sigma}m^{2}\dfrac{\delta^{(4)}(z_{i}-z^{\prime})}{\sqrt{-g(z_{i})}}D(z^{\prime},z)\bigg\}G[\{\phi(z_{j})\}]\bigg\rangle.$ (59)
@@
Let us now evaluate integrations in $y$ and $z_{i}$. First, we perform the $y$-integral which is trivial due to the presence of the delta function $\delta^{(4)}(y-z^{\prime})$. Performing $y$-integral on the delta function will simply replace the coordinate $y$ by $z^{\prime}$ in the above expression. Similarly, we can perform the $z_{i}$-integral on another delta function $\delta^{(4)}(z_{i}-z^{\prime})$. Since the coordinates $z_{i}$ and $z^{\prime}$ are independent, we move the $z_{i}$-integral inside the derivative operator $\partial^{\prime}_{\alpha}$, which acts on the delta function in the above expression, and then perform the integration. We obtain
@@
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},k)=\int d^{4}z\sqrt{-g(z)}\int d^{4}z^{\prime}\sqrt{-g(z^{\prime})}\mathcal{P}^{\mu\nu\alpha\beta}\mathcal{P}_{\alpha\beta}^{\rho\sigma}f^{*}_{\mu\nu}(z^{\prime},k)\\ \cdot\sum_{i=1}^{n}\frac{\kappa_{i}}{2}\bigg\{2\partial^{\prime}_{\rho}g^{*}_{p_{i}}(z^{\prime})\partial^{\prime}_{\sigma}D(z^{\prime},z)-g_{\rho\sigma}g^{\tau\lambda}\partial^{\prime}_{\tau}g^{*}_{p_{i}}(z^{\prime})\partial^{\prime}_{\lambda}D(z^{\prime},z)+g_{\rho\sigma}m^{2}g^{*}_{p_{i}}(z^{\prime})D(z^{\prime},z)\bigg\}\\ \cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}\sqrt{-g(z_{j})}g^{*}_{p_{j}}(z_{j})(-i)(\nabla^{2}_{z_{j}}-m^{2})\big\langle G[\{\phi(z_{j})\}]\big\rangle,$ (60)
@@
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}\frac{\kappa_{i}}{2}\int d^{4}zd^{4}z^{\prime}\left(1-\dfrac{z^{2}}{l^{2}}-\dfrac{z^{\prime 2}}{l^{2}}\right)f^{*\rho\sigma}(z^{\prime},k)\bigg\{2\partial^{\prime}_{\rho}g^{*}_{p_{i}}(z^{\prime})\partial^{\prime}_{\sigma}D(z^{\prime},z)\bigg\}\\ \cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}\sqrt{-g(z_{j})}g^{*}_{p_{j}}(z_{j})(-i)(\nabla^{2}_{z_{j}}-m^{2})\big\langle G[\{\phi(z_{j})\}]\big\rangle,$ (61)
where the traceless condition $g_{\rho\sigma}f^{*\rho\sigma}=0$ is used.
This is the main result from which we compute the leading, subleading, and sub-subleading corrections to the flat space soft factor in the following sections. Substituting all the essential elements, i.e., scalar field modes, graviton modes, and the scalar field propagator, in the above expression, keeping terms up to $\mathcal{O}(l^{-2})$, and taking the soft limit ($k\to 0$ or $\omega\to 0$) produces the final expression that can be put in the following form:
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},\omega\hat{k})=\bigg(A^{(0)}(\{p_{i}\},\omega\hat{k})+\dfrac{1}{l}A^{(1)}(\{p_{i}\},\omega\hat{k})+\dfrac{1}{l^{2}}A^{(2)}(\{p_{i}\},\omega\hat{k})\bigg)\\ \cdot\bigg(\Gamma_{n}^{(0)}(\{p_{1},...,p_{n}\})+\dfrac{1}{l}\Gamma_{n}^{(1)}(\{p_{1},...,p_{n}\})+\dfrac{1}{l^{2}}\Gamma_{n}^{(2)}(\{p_{1},...,p_{n}\})\bigg),$ (62)